Mars has a radius of about 0.38 Earth radius and a mass of only 0.11 Earth mass. Estimate the acceleration due to gravity on Mars.
1
Expert's answer
2018-09-06T16:38:08-0400
For Earth:
mg=mGM/r^2
g=GM/r^2
For Mars:
g_M=G 0.11M/(0.38r)^2 =g 0.11/(0.38)^2 =9.8 0.11/(0.38)^2 =7.5 m/s^2 .